OpenStudy (anonymous):

Which of the following is the equation of the circle with center (4, 0) and r =sqrt 5 x2 + (y - 4)2 = 5 x2 + (y + 4)2 = (x + 4)2 + y2 =

OpenStudy (anonymous):

for center (a, b) and r =c (x-a)2 + (y - b)2 = c^2